OpenStudy (anonymous):

Test to see if the series \[\sum_{n=1}^{\infty}{\frac{(-1)^{n+1}}{^4\sqrt{n}}}\]is absolutely convergent, conditionally convergent, or divergent.

OpenStudy (zarkon):

use the divergence test

OpenStudy (anonymous):

...

OpenStudy (zarkon):

\[\lim_{n\to\infty}\frac{2^n}{\sqrt[4]{n}}=\infty\]

OpenStudy (anonymous):

oops that's the wrong series, the first one is actually correct. sorry

OpenStudy (anonymous):

sorry I feel sleepy I've been doing homework for the last 6 hours.

OpenStudy (zarkon):

the first one converges conditionally

OpenStudy (anonymous):

yeah since 1/root(n,4) does not converge, but it passes the alternating series test. thanks.

OpenStudy (anonymous):

with the absolute value, it's a p series wiht p < 1 so it's divergent.